# Accurate Estimation of Mutual Information in High Dimensional Data

### v1

@@ A.1.1 InfoNCE and SMILE
Two widely used neural estimators of mutual information—InfoNCE and SMILE—can be derived from the Donsker–Varadhan (DV) representation [38] of the Kullback–Leibler (KL) divergence.
@@
Applying the DV representation to Eq. (4), one obtains:
@@
where $T(x,y)$ is a learned critic function approximating the log-density ratio. This is the MINE estimator [42, 35], which exhibits large variance empirically, specifically because of the second term.
To stabilize this estimator, SMILE [36] clips the critic before exponentiation to reduce the influence of outliers on the normalization term:
@@
Using the conditional factorization of MI in Eq. (5), one can again apply the DV representation, now to $D_{\mathrm{KL}}(p(y|x)\|p(y))$:
@@
Plugging into Eq. (5) and swapping the order of the integrals gives:
@@
which is the first step towards the InfoNCE estimator[37].
@@ A.1.2 Bilinear Critics, Gaussian Variables, and $I_{\rm CCA}$
It is well known that, for jointly Gaussian variables, MI between two random variables can be written in terms of their nonzero canonical correlations, subject to keeping enough canonical pairs [59, 46, 48]
@@ A.2 Probabilistic Critics: Variational Symmetric Information Bottleneck (VSIB)
The Variational Symmetric Information Bottleneck (VSIB) [^2] This is an instance of the more general, Deep Multivariate Information Bottleneck Framework [60, 19]. In this framework, one specifies a compression/encoder graph that is traded off against a generative/decoder graph. Each graph is then transformed into an information bound that can be optimized. can formalize MI estimation as a form of probabilistic dimensionality reduction. It introduces a latent representation for each variable—$Z_{X}$ and $Z_{Y}$—produced by separate stochastic encoders from $X$ and $Y$, respectively. Mutual information is then estimated between these latent representations using any neural MI estimator of choice.
This leads to the following objective [19]:
@@
$\approx\frac{1}{2N}\sum_{i=1}^{N}\left[\text{Tr}({\Sigma_{Z_{X}}(x_{i})})+||\vec{\mu}_{Z_{X}}(x_{i})||^{2}-k_{Z_{X}}-\ln\det(\Sigma_{Z_{X}}(x_{i}))\right],$ (25)
@@
A.3 The max-test heuristic
To define the max-test stopping heuristic, note both the training and the test data are sampled from the same $p(x,y)$. In practice, expectations for every estimator EST, such as in Eq. (7, 8, 11), are implemented with empirical sampling. That is, we form empirical densities $\pi_{\rm train}=\frac{1}{N_{\rm train}}\sum_{i=1}^{N_{\rm train}}\delta(x-x_{i},y-y_{i})$, and similar for $\pi_{\rm test}$ for the test data. Then, for example, the MINE estimator is implemented as,
@@ A.3 The max-test heuristic
    After each such step, freeze $T^{*}_{\rm train}$, and evaluate both $I_{\text{EST, train}}$ and $I_{\text{EST, test|train}}$ for whichever estimator EST is being used.
@@
    Select $\hat{T}=T^{*}_{\rm train}$ corresponding to the cycle where $I_{\text{EST, test|train}}$ is maximal.
@@
Here we justify this heuristic, which may appear to conflict standard machine learning practice, which typically report test results. However, we argue that in the case of neural MI estimators, our approach is correct. Fundamentally, this is because MI is not a linear functional of the underlying data distribution, unlike many other expectation values, which are linear in the measure. Therefore, unbiased estimates of the distribution do not yield unbiased MI estimates [61]. Because of this, it is well known, for example, that resampling approaches, such as bootstrap and cross-validation, result in errors in biases in MI estimation [40]. Here we argue that similar results extend to the current case. The key insight is that MI estimation is an estimation of a functional not a prediction task. An estimator is trained to estimate this functional with small bias on the training set, and its performance on the test set does not necessarily approximate the functional. We argue this in a few ways.
@@
The DEMINE and meta-DEMINE MI estimators [62] also attempt to produce more efficient estimators by splitting data into validation and training data sets. The training data sets can be much larger and even use task augmentation to allow the critic to be well-learned. The data efficiency of estimating information on the validation set is decoupled from learning the critic and thus can lead to better efficiency for estimating information on the validation set. The problem with this procedure is the assumption that the training data-set is large enough for the critic to be well learned and provides no way to determine when the critic is well learned.
@@
Empirically, in all cases we showed in the paper and many others, we observe that when reporting the test MI estimate, the values are consistently biased downward (cf. Fig. 3). In contrast, reporting the train MI estimate, evaluated using the best-performing model on the test set (i.e., the model checkpoint that achieved the highest test MI before overfitting), provides a more accurate estimate of the true MI in controlled synthetic setups (e.g., cf. Fig. 4).
@@ A.4 Detailed Guidelines for Estimating MI
Critic architecture should match the data modality. We use MLPs for simplicity and consistency across experiments, but CNNs or transformers [63] may be more appropriate in other tasks. In practice, any architecture that effectively captures the statistics of the input may be used, as the critic is simply an embedding-to-MI pipeline. We have verified that CNNs also work in our setup on Noisy MNIST image data from Sec. 4.3 (not shown).
@@
3. Subsampling and Max-Test Heuristic.
Start with the separable critic embedding dimension of $k_{Z}=1$. Subsample the dataset into $\gamma$ non-overlapping subsets and evaluate the estimator on each subset. Start with $\gamma=1$ (i.e., the full dataset) and compute the MI using the max-test heuristic (Fig. 3). We typically train up to 100 epochs with an additional early stopping rule if test performance stops improving for 50 epochs. More generally, the criteria should be that the training is long enough to notice saturation or decline of the test curve. Then, for $\gamma=2$, split the dataset into two halves and compute MI on each. Continue increasing $\gamma$ up to 10 (one tenth of the full dataset in each subset). Since datasets often cannot be evenly divided, allocate $\lfloor N/\gamma\rfloor$ samples to each of the first $\gamma-1$ subsets and assign the remainder to the final subset.
@@
Traditional best practice [40] suggests using inverse variance across all subsets at a given data fraction as weights in WLS, with the variance smoothed via an OLS model fit to the subset variances across all fractions. This approach assumes that the estimator variance scales as $1/N$ (i.e., $\propto\gamma$), as seen with traditional estimators like $I_{\text{KSG}}$. However, this does not hold for neural MI estimators. We have verified that these estimators exhibit relatively uniform variance across $\gamma$ (not shown), which is likely dominated not by sample size but by stochasticity in optimization, random initializations, and inherent signal structure [36]. Thus, we instead suggest assigning weights to each subset proportional to its $1/\gamma$. For $\gamma=1$, the single dataset is then given full weight; for $\gamma=2$, each of the two halves is weighted by $1/2$, and so on.
@@
Use the intercept of the linear WLS fit of $\bar{I}^{k^{*}_{Z}}_{\text{EST}}(\tilde{\gamma})$ vs $\tilde{\gamma}$ dependence to estimate MI in the $\tilde{\gamma}=0$ limit: $\hat{I}=c$. This provides an estimate at the infinite data limit and follows a procedure similar to Strong et al [6]. Alternatively, one can interpolate to $\gamma=1$ to find an MI for the full dataset size, less susceptible to statistical fluctuations. The corresponding prediction interval $\Delta I$ from the linear fit quantifies uncertainty in either estimate.
@@ A.5 Estimating Sample Size Required for Mutual Information Estimation in the Latent Variable Model
To make sense of the sample size $N$ that allows MI estimation in Fig. 5, it is helpful to recall the spiked covariance model from random matrix theory [54]. There, the data covariance matrix has the structure $\Sigma=\sigma^{2}_{n}(\mathbb{I}+\theta v^{\top}v)$, where $\sigma_{n}^{2}$ is the per-coordinate noise variance, and $v$ is a low-rank matrix with orthonormal columns, corresponding to independent signals on the background of Gaussian noise. Then the spike signal of magnitude $\theta$ separates from the sampling-induced spurious background correlations when $\theta>\sqrt{k/N}$, where $k$ is the observed dimensionality [53]. The overlap of the largest eigenvector of the data covariance matrix with the spike vectors $v$ grows continuously beyond this threshold.
@@ A.6 Implementation Details
To simplify comparisons, for all synthetic experiments (i.e., all figures except Fig. 7), we use feedforward multi-layer perceptrons (MLPs) with two hidden layers of 256 units each, initialized using Xavier uniform initialization [64], and using leaky ReLU activations. For the MNIST dataset, due to the added difficulty of the task, we use deeper networks with four hidden layers of width 512. We also used CNNs for this dataset with similar results (not shown).
@@
All unspecified hyperparameters use their default values as implemented in the PyTorch v2.0.1, SciPy v1.11.1, cca_zoo [65] v2.3.11, and statsmodels v0.14.2 libraries.
@@
For $I_{\text{CCA}}$ in Fig. 2, computing MI at a single MI level and embedding dimension $k_{Z}$ took approximately 70 seconds. For $I_{\text{KSG}}$ (as used in SI Fig. 10), the first panel took approximately 50 seconds for the entire sweep; the second panel, which had more samples, took around 500 seconds. Both were run on CPU nodes.
Overall, the total compute time, including exploratory and failed runs that led to this work, is estimated on the order of 300 compute hours across CPUs and GPUs.
Smoothing and the Max Test Heuristic.
@@
While it is possible to evaluate $I_{\text{EST, train}}$ over the full training dataset, we found that evaluating it on a representative batch is sufficient for early stopping and reporting. Larger-scale averaging can be performed at evaluation time if memory allows.
A.7 Additional Figures
Figure 8: Comparison of MI estimators in the low-dimensional, infinite-data regime. This figure replicates the setup of Fig. 1, but includes a broader selection of estimators implemented using adaptations of the code in [36, 35], with the same convention for the estimator names. The ground-truth MI is varied in discrete steps, and all estimators are run with the same batch size and sample schedule. Among the various tested methods, SMILE shows relatively low bias and captures the information levels accurately, albeit with higher variance. InfoNCE, in contrast, is consistently biased downward, especially at large information, but displays low variance. Other estimators either suffer from instabilities during training or fail to consistently track the true MI values across all cases. This comparison motivates our focus on SMILE and InfoNCE in the main experiments: they are the only methods that behave reliably across information levels and different data transformations, despite exhibiting the aforementioned trade-offs in variance and bias.

### v2

@@ A.1.1 InfoNCE and SMILE
Two widely used neural estimators of mutual information—InfoNCE and SMILE—can be derived from the Donsker–Varadhan (DV) representation (Donsker & Varadhan, 1983) of the Kullback–Leibler (KL) divergence.
@@
Applying the DV representation to Eq. equation 4, one obtains:
@@
where $T(x,y)$ is a learned critic function approximating the log-density ratio. This is the MINE estimator (Belghazi et al., 2018; Poole et al., 2019), which exhibits large variance empirically, specifically because of the second term.
To stabilize this estimator, SMILE (Song & Ermon, 2019) clips the critic before exponentiation to reduce the influence of outliers on the normalization term:
@@
Using the conditional factorization of MI in Eq. equation 5, one can again apply the DV representation, now to $D_{\mathrm{KL}}(p(y|x)\|p(y))$:
@@
Plugging into Eq. equation 5 and swapping the order of the integrals gives:
@@
which is the first step towards the InfoNCE estimator (van den Oord et al., 2018).
@@ A.1.2 Bilinear Critics, Gaussian Variables, and $I_{\rm CCA}$
It is well known that, for jointly Gaussian variables, MI between two random variables can be written in terms of their nonzero canonical correlations, subject to keeping enough canonical pairs (Kullback, 1959; Gelfand & Iaglom, 1959; Huffmann & Mittelbach, 2022):
@@ A.2 Probabilistic Critics: Variational Symmetric Information Bottleneck (VSIB)
[…]
This leads to the following objective (Abdelaleem et al., 2025):
@@
[…]
@@
A.3 The early stopping heuristic
To define the stopping heuristic, note both the training and the test data are sampled from the same $p(x,y)$. In practice, expectations for every estimator EST, such as in Eq. (7, 8, 11), are implemented with empirical sampling. That is, we form empirical densities $\pi_{\rm train}=\frac{1}{N_{\rm train}}\sum_{i=1}^{N_{\rm train}}\delta(x-x_{i},y-y_{i})$, and similar for $\pi_{\rm test}$ for the test data. Then, for example, the MINE estimator is implemented as,
@@ A.3 The early stopping heuristic
[…]
@@
[…]
@@
Here we justify this heuristic, which may appear to conflict standard machine learning practice, which typically report test results. However, we argue that in the case of neural MI estimators, our approach is correct. Fundamentally, this is because MI is not a linear functional of the underlying data distribution, unlike many other expectation values, which are linear in the measure. Therefore, unbiased estimates of the distribution do not yield unbiased MI estimates (Nemenman et al., 2001). Because of this, it is well known, for example, that resampling approaches, such as bootstrap and cross-validation, result in errors in biases in MI estimation (Holmes & Nemenman, 2019). Here we argue that similar results extend to the current case. The key insight is that MI estimation is an estimation of a functional not a prediction task. An estimator is trained to estimate this functional with small bias on the training set, and its performance on the test set does not necessarily approximate the functional. We argue this in a few ways.
@@
The DEMINE and meta-DEMINE MI estimators (Lin et al., 2019) also attempt to produce more efficient estimators by splitting data into validation and training data sets. The training data sets can be much larger and even use task augmentation to allow the critic to be well-learned. The data efficiency of estimating information on the validation set is decoupled from learning the critic and thus can lead to better efficiency for estimating information on the validation set. The problem with this procedure is the assumption that the training dataset is large enough for the critic to be well learned and provides no way to determine when the critic is well learned.
@@
Empirically, in all cases we reported in the paper and many others we tried but not reported, we observe that the test MI estimate is consistently biased downward (cf. Fig. 3). In contrast, the train MI estimate, evaluated using the best-performing model on the test set (i.e., the model checkpoint that achieved the highest test MI before overfitting), provides a more accurate estimate of the true MI in controlled synthetic setups (e.g., cf. Fig. 4).
@@ A.4 Detailed Guidelines for Estimating MI
Critic architecture should match the data modality. We use MLPs for simplicity and consistency across experiments, but CNNs or transformers (Vaswani et al., 2017) may be more appropriate in other tasks. In practice, any architecture that effectively captures the statistics of the input may be used, as the critic is simply an embedding-to-MI pipeline. We have verified that CNNs also work in our setup on Noisy MNIST image data from Sec. 4.3 (not shown).
@@
3. Subsampling and Max-Test Stopping Heuristic.
Start with the separable critic embedding dimension of $k_{Z}=1$. Subsample the dataset into $\gamma$ non-overlapping subsets and evaluate the estimator on each subset. Start with $\gamma=1$ (i.e., the full dataset) and compute the MI using the max-test stopping heuristic (Fig. 3). We typically train up to 100 epochs with an additional early stopping rule if test performance stops improving for 50 epochs. More generally, the criteria should be that the training is long enough to notice saturation or decline of the test curve. Then, for $\gamma=2$, split the dataset into two halves and compute MI on each. Continue increasing $\gamma$ up to 10 (one tenth of the full dataset in each subset). Since datasets often cannot be evenly divided, allocate $\lfloor N/\gamma\rfloor$ samples to each of the first $\gamma-1$ subsets and assign the remainder to the final subset.
@@
Traditional best practice (Holmes & Nemenman, 2019) suggests using inverse variance across all subsets at a given data fraction as weights in WLS, with the variance smoothed via an OLS model fit to the subset variances across all fractions. This approach assumes that the estimator variance scales as $1/N$ (i.e., $\propto\gamma$), as seen with traditional estimators like $I_{\text{KSG}}$. However, this does not hold for neural MI estimators. We have verified that these estimators exhibit relatively uniform variance across $\gamma$ (not shown), which is likely dominated not by sample size but by stochasticity in optimization, random initializations, and inherent signal structure (Song & Ermon, 2019). Thus, we instead suggest assigning weights to each subset proportional to its $1/\gamma$. For $\gamma=1$, the single dataset is then given full weight; for $\gamma=2$, each of the two halves is weighted by $1/2$, and so on.
@@
[…]
@@ A.5 Estimating Sample Size Required for Mutual Information Estimation in the Latent Variable Model
To make sense of the sample size NN that allows MI estimation in Fig. 5, it is helpful to recall the spiked covariance model from random matrix theory (Potters & Bouchaud, 2020). There, the data covariance matrix has the structure $\Sigma=\sigma^{2}_{n}(\mathbb{I}+\theta v^{\top}v)$, where $\sigma_{n}^{2}$ is the per-coordinate noise variance, and $v$ is a low-rank matrix with orthonormal columns, corresponding to independent signals on the background of Gaussian noise. Then the spike signal of magnitude $\theta$ separates from the sampling-induced spurious background correlations when $\theta>\sqrt{k/N}$, where $k$ is the observed dimensionality (Baik et al., 2005). The overlap of the largest eigenvector of the data covariance matrix with the spike vectors $v$ grows continuously beyond this threshold.
@@ A.6 Implementation Details
To simplify comparisons, for all synthetic experiments (i.e., all figures except Fig. 7), we use feedforward multi-layer perceptrons (MLPs) with two hidden layers of 256 units each, initialized using Xavier uniform initialization (Glorot & Bengio, 2010), and using leaky ReLU activations. For the MNIST dataset, due to the added difficulty of the task, we use deeper networks with four hidden layers of width 512. We also used CNNs for this dataset with similar results (not shown).
@@
All unspecified hyperparameters use their default values as implemented in the PyTorch v2.0.1, SciPy v1.11.1, cca_zoo (Chapman & Wang, 2021) v2.3.11, and statsmodels v0.14.2 libraries.
@@
For $I_{\text{CCA}}$ in Fig. 2, computing MI at a single MI level and embedding dimension $k_{Z}$ took approximately 70 seconds. In Appx. A.7.2, we benchmark our approach against other methods; so here we report their compute requirements as well. For $I_{\text{KSG}}$ (as used in Appx. Fig. 9), the first panel took approximately 50 seconds for the entire sweep; the second panel, which had more samples, took around 500 seconds. Both were run on CPU nodes. $I_{\text{LMI}}$ took about 5 minutes to run on a desktop CPU. $I_{\text{Geodesic}}$ took 2.25 days to run on the CPU nodes of the cluster and needed more than $128$ GB of RAM to complete. Neither of these estimates sample-size dependent biases or confidence intervals. That is, our approach compares favorably to others in terms of computing needs.
Overall, the total compute time, including exploratory and failed runs that led to this work, is estimated on the order of 500 compute hours across CPUs and GPUs.
Smoothing and the Stopping Heuristic.
@@
While it is possible to evaluate $I_{\text{EST, train}}$ over the full training dataset, we found that evaluating it on a representative batch is sufficient for early stopping and reporting. Larger-scale averaging can be performed at evaluation time if memory allows. However, for completeness, we performed such an evaluation over the full training dataset for the task mentioned later in Sec. 2 and the performance is consistent, albeit with technical caveats to prevent memory related crashes.
A.7 Evaluations with Other Estimators and Datasets
In this section, we provide additional evaluations of our pipeline beyond the experiments reported in the main text. Specifically, we study: (i) alternative estimators in the infinite-data, low-dimensional regime, which motivated our choice of InfoNCE and SMILE; (ii) other popular estimators, including KSG (Kraskov et al., 2004), Sliced MI (SMI) (Goldfeld & Greenewald, 2021), Geodesic MI (Marx & Fischer, 2022), and LMI (Gowri et al., 2024), in the finite-data regime of our teacher network task (Fig. 4); and (iii) evaluations on the benchmark datasets introduced in Czyz et al. (2023), which allows us to compare the traditional use of InfoNCE with its modification in our pipeline.
A.7.1 Infinite Data Regime
Figure 8: Comparison of MI estimators in the low-dimensional, infinite-data regime. We replicate the setup of Fig. 1 but include a broader selection of estimators (implementations adapted from Song & Ermon (2019); Poole et al. (2019)). True MI is varied in discrete steps, and all estimators are run with identical batch size and sample schedules. SMILE tracks the ground truth most closely but with higher variance, while InfoNCE is biased downward at high MI yet remains stable. Other estimators either diverge during training or fail to capture the correct trends. This motivates focusing on SMILE and InfoNCE in our main study.
A.7.2 Finite Data Regime
KSG.
The KSG estimator (Kraskov et al., 2004) with code adapted from Czyz et al. (2023), was used to estimate information contained in $K=500$-dimensional data generated by teacher networks with latent dimensionality $K_{Z}=10$ and true MI of 4 bits, as used in Fig. 6 for neural estimators. Best practices, were used to estimate the MI with KSG estimator (Holmes & Nemenman, 2019) at different numbers of nearest neighbors $k$. Specifically, as for NN estimators, data was partitioned into $\gamma$ non-overlapping partitions and estimates for each partition were found. The mean information and standard deviation using $\gamma$ partitions were plotted versus the number of partitions at different values of $k$ in Fig. 9. The KSG estimator does not asymptotically approach the correct value of information at $\gamma=0$, corresponding to the infinite data limit, and linear extrapolations (done the same way as in Fig. 6) is unreliable for KSG in high dimensions. Best practice is to not use the estimator in this case (Holmes & Nemenman, 2019).
Figure 9: Failure of the KSG estimator in high dimensions. Estimation of $I(X;Z)$ with KSG (Kraskov et al., 2004) on $K=500$-dimensional teacher data with latent dimension $K_{Z}=10$ and ground-truth MI of 4 bits. Left: undersampled ($N=256$) regime; Right: well-sampled ($N=2^{14}$). For different nearest-neighbor values $k$, MI estimates remain far below ground truth (1–2.8 bits) and do not converge under extrapolation. In contrast, neural estimators recover near-perfect values (Fig. 6). This illustrates that KSG is unreliable for high-dimensional data, consistent with Holmes & Nemenman (2019).
Sliced MI.
Sliced Mutual Information (SMI) (Goldfeld & Greenewald, 2021) does not estimate true MI but a surrogate quantity that violates the data processing inequality. While SMI may be useful for representation learning, comparing it directly to MI estimators is misleading. On our high-dimensional teacher network benchmark, available implementations yielded values far below ground truth (e.g., 1.13 bits vs. 4 bits), highlighting its inapplicability as a general MI estimator.
Geodesic MI.
Geodesic-MI (Marx & Fischer, 2022) proved computationally prohibitive: runs required multiple days and >128>128GB RAM. For the better-sampled case ($N=2^{14}$), it produced estimates of 1.86 and 1.72 bits (for $k=3,10$), far below the 4-bit ground truth, showing both poor scalability and low accuracy compared to our approach.
LMI.
Latent MI (LMI) estimators (Gowri et al., 2024) assume linear-Gaussian structure and are unreliable in nonlinear settings. On our teacher network task (true MI 4 bits), LMI gave 0.65 bits (undersampled) and 1.6 bits (better-sampled), versus our $4.1\pm 0.2$ and $3.7\pm 0.1$. Unlike LMI’s two-step compression–estimation procedure, our end-to-end training captures nonlinear dependencies directly.
A.7.3 Other Benchmark Datasets
Czyz et al. (2023) introduced a diverse suite of datasets ($1\leq K\leq 50$) with analytically tractable MI values. They reported that InfoNCE generally outperformed alternatives (Figure 2 in their paper). We revisit these datasets (10,000 samples) using the same critic architecture as in their study (a concatenated critic with two hidden layers of width 16), but apply our bias correction, error bar estimation pipeline with held-out testing (9k train, 1k test samples).
Table 2 compares ground-truth MI, results from Czyz et al. (2023), and our pipeline. Grey rows mark unreliable fits ($\delta>0.1$ or $\gamma_{\max}\leq 5$). Our method consistently matches or exceeds baseline InfoNCE performance while uniquely providing confidence intervals.
For challenging datasets such as spirals, performance improves substantially with more expressive critics (hidden layers of width 256). Table 3 illustrates this case. Overall, our pipeline not only matches or surpasses state-of-the-art results, but is the only one offering error bars and internal consistency checks.
Table 2: Benchmark results on datasets from Czyz et al. (2023). True MI, results reported by Czyz et al. (2023) using simple InfoNCE implementation, and our pipeline with error bars. Grey rows indicate unreliable fits ($\delta>0.1$ or $\gamma_{\max}<5$). Our method consistently achieves estimates similar to or better than (reported in boldface) the rivals, while uniquely reporting confidence intervals. Results reported in nats to match Czyz et al. (2023).
Task True MI simple InfoNCE Ours $\gamma_{\textrm{max}}$ $\delta$
Uniform 1 × 1 (additive noise=0.1) 1.71 1.7 1.67 ± 0.03 10 0.041
Uniform 1 × 1 (additive noise=0.75) 0.33 0.3 0.33 ± 0.02 10 0.026
Bimodal 1 × 1 0.41 0.4 0.39 ± 0.02 5 0.195
Bivariate normal 1 × 1 0.41 0.4 0.39 ± 0.03 9 0.087
Asinh @ Student-t 1 × 1 (dof=1) 0.22 0.2 0.24 ± 0.02 9 0.064
Asinh @ Student-t 2 × 2 (dof=1) 0.43 0.4 0.40 ± 0.02 9 0.068
Asinh @ Student-t 3 × 3 (dof=2) 0.29 0.2 0.26 ± 0.03 10 0.098
Asinh @ Student-t 5 × 5 (dof=2) 0.45 0.3 0.34 ± 0.07 10 0.058
Half-cube @ Bivariate normal 1 × 1 0.41 0.4 0.39 ± 0.01 10 0.077
Half-cube @ Multinormal 25 × 25 (2-pair) 1.02 0.8 1.02 ± 0.07 5 0.234
Half-cube @ Multinormal 3 × 3 (2-pair) 1.02 1.0 1.00 ± 0.02 10 0.083
Half-cube @ Multinormal 5 × 5 (2-pair) 1.02 1.0 1.01 ± 0.03 9 0.062
Multinormal 2 × 2 (dense) 0.29 0.3 0.30 ± 0.02 10 0.086
Multinormal 25 × 25 (dense) 1.29 1.2 1.26 ± 0.03 10 0.03
Multinormal 3 × 3 (dense) 0.41 0.4 0.40 ± 0.02 10 0.093
Multinormal 5 × 5 (dense) 0.59 0.6 0.60 ± 0.03 10 0.098
Multinormal 50 × 50 (dense) 1.62 1.4 1.60 ± 0.03 10 0.062
Multinormal 2 × 2 (2-pair) 1.02 1.0 1.02 ± 0.03 10 0.059
Multinormal 25 × 25 (2-pair) 1.02 0.9 0.98 ± 0.05 5 0.162
Multinormal 3 × 3 (2-pair) 1.02 1.0 1.00 ± 0.02 5 0.166
Multinormal 5 × 5 (2-pair) 1.02 1.0 1.02 ± 0.02 7 0.045
Normal CDF @ Bivariate normal 1 × 1 0.41 0.4 0.37 ± 0.02 10 0.034
Normal CDF @ Multinormal 25 × 25 (2-pair) 1.02 0.8 0.94 ± 0.12 9 0.009
Normal CDF @ Multinormal 3 × 3 (2-pair) 1.02 0.9 0.92 ± 0.04 10 0.048
Normal CDF @ Multinormal 5 × 5 (2-pair) 1.02 0.9 0.93 ± 0.04 10 0.046
Spiral @ Multinormal 25 × 25 (2-pair) 1.02 0.7 0.97 ± 0.24 9 0.022
Spiral @ Multinormal 3 × 3 (2-pair) 1.02 0.6 0.69 ± 0.04 10 0.037
Spiral @ Multinormal 5 × 5 (2-pair) 1.02 0.6 0.64 ± 0.04 10 0.035
Spiral @ Normal CDF @ Multinormal 25 × 25 (2-pair) 1.02 0.8 0.89 ± 0.10 6 0.096
Spiral @ Normal CDF @ Multinormal 3 × 3 (2-pair) 1.02 0.9 0.89 ± 0.03 10 0.042
Spiral @ Normal CDF @ Multinormal 5 × 5 (2-pair) 1.02 0.9 0.88 ± 0.04 9 0.064
Student-t 1 × 1 (dof=1) 0.22 0.1 0.21 ± 0.03 5 0.195
Student-t 2 × 2 (dof=1) 0.43 0.3 0.27 ± 0.09 10 0.071
Student-t 2 × 2 (dof=2) 0.19 0.2 0.20 ± 0.02 9 0.054
Student-t 3 × 3 (dof=2) 0.29 0.2 0.26 ± 0.04 10 0.031
Student-t 3 × 3 (dof=3) 0.18 0.1 0.18 ± 0.03 9 0.021
Student-t 5 × 5 (dof=2) 0.45 0.4 0.37 ± 0.06 10 0.052
Student-t 5 × 5 (dof=3) 0.3 0.2 0.23 ± 0.04 10 0.054
Swiss roll 2 × 1 0.41 0.4 0.38 ± 0.02 6 0.048
Wiggly @ Bivariate normal 1 × 1 0.41 0.4 0.39 ± 0.01 5 0.184
Table 3: Improved results on spiral datasets with larger critics. For challenging cases where smaller critics underperform, increasing the hidden layer width from 16 to 256 substantially improves accuracy. Grey rows indicate unreliable fits.
Task True MI simple InfoNCE Ours (16 hidden) Ours (256 hidden) $\gamma_{\textrm{max}}$ $\delta$
Multinormal 25 × 25 1.02 0.7 0.97 ± 0.24 1.11 ± 0.17 10 0.076
Multinormal 3 × 3 1.02 0.6 0.69 ± 0.04 0.99 ± 0.05 10 0.061
Multinormal 5 × 5 1.02 0.6 0.64 ± 0.04 1.04 ± 0.08 5 0.136
Normal CDF @ Multinormal 25 × 25 1.02 0.8 0.89 ± 0.10 0.93 ± 0.10 10 0.067
Normal CDF @ Multinormal 3 × 3 1.02 0.9 0.89 ± 0.03 1.00 ± 0.04 10 0.005
Normal CDF @ Multinormal 5 × 5 1.02 0.9 0.88 ± 0.04 0.99 ± 0.03 10 0.076
A.8 Additional Figures